# Gamma: Toward Generic Image Assessment with Mixture of Assessment Experts

### v1

@@ 4.1 Datasets
Table 3: The impact of different number of experts in adaptive experts. We use naive prompt strategy for ablation.
Dataset LIVEC CSIQ TID2013 AGIQA3k UWIQA
Experts Number SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Zero Expert 0.765 0.792 0.852 0.898 0.792 0.826 0.800 0.866 0.750 0.768
One Expert 0.842 0.866 0.945 0.963 0.918 0.931 0.866 0.908 0.859 0.873
Three Experts 0.851 0.871 0.949 0.966 0.926 0.934 0.870 0.910 0.863 0.878
Five Experts 0.854 0.889 0.951 0.965 0.926 0.934 0.872 0.911 0.860 0.876
4.2 Implementation Details
Following the settings in (Su et al., 2020; Ke et al., 2021), we randomly divide each dataset into 80% for training and 20% for testing. The training dataset is a mixture of the training sets of each dataset and we test Gamma on each test data separately. This process is repeated 10 times, and the median of the 10 scores is reported as the final score. We use the pre-trained weight of UniQA (Zhou et al., 2024b), which uses CLIP-B/16 as multimodal encoder. We freeze the CLIP visual and text encoders, training only the MoAE module and adapter. For the unified training, we train the model for 10 epochs with a batch size of 8. The initial learning rate is set to 2e-5. We normalize the MOS/DMOS scale to [0, 1] and utilize Adam optimizer (Kingma, 2014) and MSE loss to optimize the model. For the task-specific training, we use different training settings according to the task and size of datasets. More training details are provided in the appendix.
Table 4: The impact of different model configuration in the proposed MoAE module.
@@ 4.2 Implementation Details
Table 5: The impact of adding MoAE to different numbers of layers for training. Time means training time.
@@
4.3 Main Results
Our MoAE-equipped model can be used for task-specific training and mixed training, both of which can achieve state-of-the-art (SOTA) performance, as shown in Table 1.
Mixed Training. We conduct mixed training on 12 image assessment datasets. The trained model can be used to assess the images from these datasets. The experimental results are reported in Table 1. When compared with other mixed training models, such as StairIQA and PromptIQA, our method exhibits powerful and superior performance on each dataset. More importantly, our method can also be used to IAA tasks and demonstrates excellent performance. It is worth noting that our mixed training model even achieves results comparable to task-specific models on datasets such as KADID, KonIQA, SPAQ, and GFIQA. These results demonstrate that our approach can be effectively applied to different image assessment scenarios.
Comparison with other mixed training methods. We conduct a detailed comparison with other mixed training methods, including LIQE, UNIQUE and Q-Align, with the same training datasets. We also perform the cross-dataset validation with LIQE and UNIQUE. Please refer to Appendix B for datails.
4.4 Ablation Studies
We conduct detailed ablation studies to validate the effectiveness of our proposed modules. Note that we use naive prompt strategy (described in Section 3.2) to perform all ablations unless otherwise specified. We uniformly use 12 datasets for ablation experiments. Considering the page limit, we only show the datasets with relatively large differences in results. More ablations about prompt sensitivity and model efficiency please refer to Appendix C and D.
Figure 5: (a) Average performance when using different numbers of experts. Adding adaptive experts significantly improves model performance. (b) Average activations of three experts in the last layer of the visual encoder with naive prompts. Image evaluations of different scenes have different activation patterns.
Table 6: Results when only one adaptive expert is activated. The weights factors of other experts are set to 0. It can be observed that different experts focus on different datasets.
@@ 4.4 Ablation Studies
Table 7: Generalization capability validation on the exBeDDE and ECIQAD datasets. The “Pretrained weight” denotes the model weight of mixed training. We can notice that loading pretrained weight for initialization can improve model performance.
(a) Results on the exBeDDE datasets.
@@
(b) Results on the ECIQAD datasets.
@@
FRIQUEE (Ghadiyaram & Bovik, 2017) 0.663 0.656
@@
Merging features with factor $\sigma$. Table 4 also demonstrates the effect of merging features of shared and adaptive experts with factor $\sigma$. We notice that this strategy improves the model performance on different datasets, especially on AVA and AGIQA3k. These results show that it is beneficial to utilize aligned features at the beginning of training and partially using features from adaptive experts.
@@
Activation patterns of different datasets. We visualize the average activation degree of three experts in the last layer of Gamma’s visual encoder for different datasets, as shown in Figure LABEL:fig:ab-activation. We can observe that the activation patterns are different for different scenarios. Specifically, the natural image assessment datasets, e.g., LIVE, CSIQ, KADID, show different activation patterns from the face IQA dataset GFIQA and the underwater IQA dataset UWIQA. The synthetic distortion and authentic distortion dataset in nature IQA also have different activation patterns. These indicate that our MoAE module can assign experts with different activation levels to images of different scenarios, thereby capturing the discriminative features effectively.
@@
4.5 Generalization Capability Validation
We further validate the generalization capability of our method on two datasets, exBeDDE and ECIQAD. The exBeDDE is a dehazed IQA dataset and the ECIQAD is an enhanced colonoscopy IQA dataset, which belong to completely different evaluation domains compared to the used datasets in mixed training. We use naive prompt strategy for training and testing. The experimental results are reported in Table 7. We notice that our method can achieve competitive performance on these two datasets, showing the effectiveness and generalization capability of our method. More importantly, when we load the pretrained weight of Gamma for initialization, the performance of both datasets is improved and our method achieves the SOTA results. This indicates that our pretrained Gamma can be an effective foundation model to aid other assessment fields.
5 Conclusion
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A More Implementation Details
A.1 Evaluation Criteria
We use Spearman’s Rank-Order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC) as criteria to measure the performance of IQA models. Both coefficients range from 0 to 1, with higher values indicating better performance.
A.2 Training Details
Table 8: Training settings for different datasets.
@@ A.2 Training Details
CSIQ (Larson & Chandler, 2010) SDN-IQA 50 8
@@
CLIVE (Ghadiyaram & Bovik, 2015) ADN-IQA 50 8
@@
A.3 Datasets
Table 9: Detail information about the 14 used datasets.
Dataset Task Image Number Label Type Range
@@ A.3 Datasets
CSIQ (Larson & Chandler, 2010) 866 DMOS [0, 1]
@@
LIVEC (Ghadiyaram & Bovik, 2015) 1,162 MOS [1, 100]
@@
Table 10: Detail information about the 14 used datasets.
Method Trainable Parms FLOPs Inference time KonIQ SRCC KADID SRCC
Q-Align (Wu et al., 2023) 8.2B (8200M) - 0.1s 0.938 0.934
LIQE (Zhang et al., 2023) 151M 17.40G 0.02s 0.919 0.930
Gamma 122.8M 28.45G 0.025s 0.939 0.962
A.4 Details of the Scene-based Differential Prompt
Table 11: Text prompts used in Scene-based Differential Prompt.
@@ A.4 Details of the Scene-based Differential Prompt
Appendix B Comparison with other mixed training methods
Table 12: Comparison with LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021) when using the same training data.
Dataset LIVE CSIQ KADID BID LIVEC KonIQ Average
Table 13: Cross-dataset validation when using the same training data as LIQE and UNIQUE. The subscripts “$s$” and “$r$” stand for models trained on KADID and KonIQ, respectively. Gamma uses the same 6 datasets as LIQE and UNIQUE for training and validation. Gamma†† uses 12 datasets (include TID2012 and SPAQ) for training and performs zero-shot cross dataset validation on other datasets. Gamma†† significantly enhances the zero-shot performance on other datasets, showing the effectiveness of mix-training via MoAE and SDP methods.
Dataset TID2013 SPAQ AIGC2023 Average
NIQE 0.314 0.578 - 0.446
DBCNNs 0.686 0.412 0.730 0.609
PaQ2PiQ 0.423 0.823 0.643 0.630
MUSIQr 0.584 0.853 0.736 0.724
UNIQUE 0.768 0.838 0.761 0.789
LIQE 0.811 0.881 0.744 0.812
Gamma 0.805 0.894 0.770 0.823
Gamma†† - - 0.818 -
In this subsection, we conduct a more detailed comparison with other mixed training methods. We first compare with LIQE and UNIQUE using the same training data and data splitting ratios. As shown in Table 12, our method achieves better performance on most datasets than LIQE and UNIQUE, especially on the KADID (+2.5% SRCC) and BID (+2.6% SRCC) datasets compared with LIQE. On other datasets, i.e., LIVE and LIVEC, our model also achieves competitive results. Overall, our model has superior performance on these five datasets.
In addition, we conduct cross dataset validation under this setting. As shown in Table 13, our method achieves highly competitive results on TID2013 and SPAQ, demonstrating the strong generalization capability of our method. Note that using mix-training of more dataset can significantly enhance the zero-shot performance on other datasets, i.e., 0.770 SRCC to 0.818 SRCC on AIGC2023 dataset, demonstrating the effectiveness of mix-training via MoAE and SDP methods.
Compared with Q-Align, as shown in Table 14, our method achieves better results on KonIQ and KADID, and is also highly competitive on SPAQ.
Table 14: Comparison with Q-Align (Wu et al., 2023) when using the same training data.
Dataset KonIQ SPAQ KADID Average
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Q-Align 0.938 0.945 0.931 0.933 0.934 0.935 0.934 0.938
Gamma 0.940 0.950 0.928 0.932 0.962 0.964 0.943 0.949
Appendix C Sensitivity analysis of prompt
Table 15: Sensitivity analysis of prompt. Quality prompt is {bad-quality, poor-quality, fair-quality, good-quality, perfect-quality}; General prompt replaces the scene prompt (see Table 11) to “general”, e.g., {underwear bad-quality image} to {general bad-quality image}.

### v2

@@ 4.2. Implementation Details
Table 3. The impact of different number of experts in adaptive experts. We use naive prompt strategy for ablation.
Dataset LIVEC CSIQ TID2013 AGIQA3k UWIQA Average
Experts Number SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Zero Expert 0.765 0.792 0.852 0.898 0.792 0.826 0.800 0.866 0.750 0.768 0.792 0.83
One Expert 0.842 0.866 0.945 0.963 0.918 0.931 0.866 0.908 0.859 0.873 0.886 0.908
Three Experts 0.851 0.871 0.949 0.966 0.926 0.934 0.870 0.910 0.863 0.878 0.892 0.912
Five Experts 0.854 0.889 0.951 0.965 0.926 0.934 0.872 0.911 0.860 0.876 0.893 0.915
Table 4. The impact of different model configuration in the proposed MoAE module.
@@
4.3. Main Results
Our MoAE-equipped model can be used for task-specific training and mixed training, both of which can achieve state-of-the-art (SOTA) performance, as shown in Table 1. We conduct a comprehensive performance comparison of the Gamma with 30 SOTA IQA methods on the 6 IQA tasks. We compare natural scene IQA methods, including HyperIQA (Su et al., 2020), MUSIQ (Ke et al., 2021), TOPIQ (Chen et al., 2024a), DEIQT (Qin et al., 2023), LoDA (Xu et al., 2024), UniQA (Zhou et al., 2024b), UNIQUE (Zhang et al., 2021), LIQE (Zhang et al., 2023), StairIQA (Sun et al., 2023) and PromptIQA (Chen et al., 2024b); face IQA methods, including IFQA (Jo et al., 2023), StyleGAN-IQA from (Su et al., 2023a), CR-FIQA (Boutros et al., 2023), GPFIQA (Su et al., 2023a); AIGC IQA methods, including DBCNN (Zhang et al., 2018), CLIPIQA (Wang et al., 2023) and PSCR (Yuan et al., 2023); underwater IQA methods, including FDUM (Yang et al., 2021a), UIQI (Liu et al., 2023), URanker (Guo et al., 2023), UCIQE (Yang and Sowmya, 2015) and HPUIQA (Liu et al., 2023), and IAA methods, including VILA (Ke et al., 2023), TANet (He et al., 2022), MaxViT (Tu et al., 2022), TAVAR (Li et al., 2023a) and ELTA (Liu et al., 2024a).
Mixed Training. We conduct mixed training on 12 image assessment datasets. The trained model can be used to assess the images from these datasets directly. The experimental results are reported in Table 1. When compared with other mixed training models, our method exhibits competitive results on each dataset. More importantly, our method can also be used to IAA tasks and demonstrates excellent performance. It is worth noting that our mixed training model even achieves results comparable to task-specific models on datasets such as KADID, KonIQA, SPAQ, and GFIQA. These results demonstrate that our approach can be effectively applied to different image assessment scenarios.
Table 5. The impact of adding MoAE to different numbers of layers for training. Time means training time.
@@ 4.3. Main Results
4.4. Ablation Studies
We conduct detailed ablation studies to validate the effectiveness of our modules. Note that we use naive prompt strategy (described in Section 3.2) to perform all ablations unless otherwise specified. We uniformly use 12 datasets for ablation experiments. Considering the page limit, we only show the datasets with relatively large differences in results. More ablations about prompt sensitivity and model efficiency please refer to Supplementary Materials.
Table 6. Results when only one adaptive expert is activated. The weights factors of other experts are set to 0. It can be observed that different experts focus on different datasets.
@@ 4.4. Ablation Studies
Table 7. Comparison with LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021) when using the same training data.
Dataset LIVE CSIQ KADID BID LIVEC KonIQ Average
Table 8. Comparison with Q-Align (Wu et al., 2023) when using the same training data.
Dataset KonIQ SPAQ KADID Average
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Q-Align 0.938 0.945 0.931 0.933 0.934 0.935 0.934 0.938
Gamma 0.940 0.950 0.928 0.932 0.962 0.964 0.943 0.949
Table 9. Generalization capability validation on the exBeDDE and ECIQAD datasets. The “Pretrained weight” denotes the model weight of mixed training. Loading pretrained weight for initialization improves model performance.
(a) Results on the exBeDDE dataset.
@@
(b) Results on the ECIQAD dataset.
@@
FRIQUEE (Ghadiyaram and Bovik, 2017) 0.663 0.656
@@
Figure 5. (a) Average performance when using different numbers of experts. Adding adaptive experts significantly improves model performance. (b) Average activations of three experts in the last layer of the visual encoder. Image evaluations of different scenes have different activation patterns.
@@
Merging features with factor $\sigma$. Table 4 also demonstrates the effect of merging features of shared and adaptive experts with factor $\sigma$. We notice that this strategy improves the model performance on different datasets, especially on AVA and AGIQA3k. These results show that it is beneficial to utilize aligned features at the beginning of training and partially using features from adaptive experts. We plot the value change of $\sigma$ during training in Figure 6. The absolute value of $\sigma$ increases gradually and eventually stabilize. This means that the adaptive expert is constantly fitting the bias between different datasets. The value is negative, suggesting that the adaptive experts remove the preference scores from the shared expert, similar to a denoising process.
@@
Activation patterns of different datasets. We visualize the average activation degree of three experts in the last layer of Gamma’s visual encoder for different datasets, as shown in Figure LABEL:fig:ab-activation. We observe that the activation patterns are different for different scenarios. Specifically, the natural IQA datasets, e.g., LIVE, CSIQ, KADID, show different activation patterns from the face IQA dataset GFIQA and the underwater IQA dataset UWIQA. The synthetic and authentic distortion datasets in natural IQA also have different activation patterns. These indicate that our MoAE module can assign experts with different activation levels to images of different scenarios, thereby capturing the discriminative features effectively.
@@
Figure 6. The change of merging factor $\sigma$ during training. The $\sigma$ is from the MoAE of the last visual encoder layer.
4.5. Comparison with Mixed Training Methods
In this subsection, we conduct a more detailed comparison with other mixed training methods. We first compare with LIQE and UNIQUE using the same training data and data splitting ratios. As shown in Table 7, our method achieves better performance on most datasets than LIQE and UNIQUE, especially on the KADID (+2.5% SRCC) and BID (+2.6% SRCC) datasets compared with LIQE. On other datasets, i.e., LIVE and LIVEC, our model also achieves competitive results. Overall, our model has superior performance on these five datasets. Compared with Q-Align, as shown in Table 8, our method achieves better results on KonIQ and KADID, and is also highly competitive on SPAQ.
4.6. Generalization Capability Validation
We further validate the generalization capability of our method on two datasets, exBeDDE and ECIQAD. The exBeDDE is a dehazed IQA dataset and the ECIQAD is an enhanced colonoscopy IQA dataset, which belong to completely different domains compared to the used datasets in mixed training. We use naive prompt strategy for training and testing. The experimental results are reported in Table 9. We notice that our method can achieve competitive performance on these two datasets, showing the effectiveness and generalization capability of our method. More importantly, when we load the pretrained weight of Gamma for initialization, the performance of both datasets is improved and our method achieves the SOTA results. This indicates that our pretrained Gamma can be an effective foundation model to aid other assessment fields.
5. Conclusion
@@ Appendix A More Implementation Details
Table 10. Training settings for different datasets.
@@
CSIQ (Larson and Chandler, 2010) SDN-IQA 50 8
@@
CLIVE (Ghadiyaram and Bovik, 2015) ADN-IQA 50 8
@@
Table 11. Detail information about the 14 used datasets.
Dataset Task Image Number MOS Type Range
@@
CSIQ (Larson and Chandler, 2010) 866 DMOS [0, 1]
@@
LIVEC (Ghadiyaram and Bovik, 2015) 1,162 MOS [1, 100]
@@
Table 12. Cross-dataset validation when using the same training data as LIQE (Zhang et al., 2023) and UNIQUE (Zhang et al., 2021). The subscripts “$s$” and “$r$” stand for models trained on KADID and KonIQ, respectively. Gamma uses the same 6 datasets as LIQE and UNIQUE for training and validation. Gamma†† uses 12 datasets (include TID2012 and SPAQ) for training and performs zero-shot cross dataset validation on other datasets. Gamma†† significantly enhances the zero-shot performance on other datasets, showing the effectiveness of mix-training via MoAE and SDP.
Dataset TID2013 SPAQ AIGC2023 Average
NIQE 0.314 0.578 - 0.446
DBCNNs 0.686 0.412 0.730 0.609
PaQ2PiQ 0.423 0.823 0.643 0.630
MUSIQr 0.584 0.853 0.736 0.724
UNIQUE 0.768 0.838 0.761 0.789
LIQE 0.811 0.881 0.744 0.812
Gamma 0.805 0.894 0.770 0.823
Gamma†† - - 0.818 -
Table 13. Generalization capability validation on the BAID (Yi et al., 2023) and ICAA17K (He et al., 2023) datasets. We use the pre-trained weight of Gamma for model initialization.
(a) Results on the BAID dataset.
Method SRCC PLCC
NIMA (Talebi and Milanfar, 2018) 0.393 0.382
MPada (Sheng et al., 2018) 0.437 0.425
MLSP (Hosu et al., 2019) 0.441 0.430
BIAA (Zhu et al., 2020) 0.389 0.376
TANet (He et al., 2022) 0.453 0.437
BAID (Yi et al., 2023) 0.473 0.467
Ours 0.496 0.556
(b) Results on the ICAA17K dataset.
Method SRCC PLCC
NIMA (Talebi and Milanfar, 2018) 0.809 0.815
MPada (Sheng et al., 2018) 0.810 0.819
MLSP (Hosu et al., 2019) 0.826 0.837
BIAA (Zhu et al., 2020) 0.820 0.829
MUSIQ (Ke et al., 2021) 0.835 0.841
TANet (He et al., 2022) 0.829 0.836
MaxViT (Tu et al., 2022) 0.855 0.874
ICAA (He et al., 2023) 0.873 0.890
Ours 0.877 0.893
Table 14. Text prompts used in Scene-based Differential Prompt.
@@
Table 15. Model complexity comparison with LIQE and Q-Align.
Method Trainable Parms FLOPs Inference time KonIQ SRCC KADID SRCC
Q-Align (Wu et al., 2023) 8.2B (8200M) - 0.1s 0.938 0.934
LIQE (Zhang et al., 2023) 151M 17.40G 0.02s 0.919 0.930
Gamma 122.8M 28.45G 0.025s 0.939 0.962
Table 16. Sensitivity analysis of prompt. Quality prompt is {bad-quality, poor-quality, fair-quality, good-quality, perfect-quality}; General prompt replaces the scene prompt (see Table 14) to “general”, e.g., {underwear bad-quality image} to {general bad-quality image}.
@@
A.1. Evaluation Criteria
We use Spearman’s Rank-Order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC) as criteria to measure the performance of IQA models. Both coefficients range from 0 to 1, with higher values indicating better performance. We randomly crop a 224x224 image patch from the original image and perform score prediction, following (Qin et al., 2023; Zhang et al., 2023). This process is repeated 10 times, and the average of the 10 predicted scores is used as the final quality score of the image.
A.2. Training Details
A.3. Datasets
A.4. Details of the Scene-based Differential Prompt
Appendix B Cross Dataset Validation
We conduct cross dataset validation to evaluate the generalization ability of our method. We compare with LIQE and UNIQUE using the same training data and data splitting ratios. As shown in Table 12, our method achieves highly competitive results on TID2013 and SPAQ, demonstrating the strong generalization capability of our method. Note that using mix-training of more dataset can significantly enhance the zero-shot performance on other datasets, i.e., 0.770 SRCC to 0.818 SRCC on AIGC2023 dataset, demonstrating the effectiveness of mix-training via MoAE and SDP methods.
Appendix C More Generalization Capability Validation
Gamma can be used as a pre-trained model and easily applied to new tasks after fine-tuning. We select the artistic IAA dataset BAID (Yi et al., 2023) and color IAA dataset ICAA17K (He et al., 2023) for further performance evaluation. BAID is a large-scale artistic image aesthetics assessment (AIAA) dataset, which consists of 60,337 artistic images covering various art forms. ICAA17K is a color-oriented IAA dataset, containing 17K images. We fine-tune Gamma for 10 epochs on them. As shown in the Table 13, Gamma achieves SOTA performance on two datasets, although it is not specifically designed for them. These results further verify the powerful generalization ability of Gamma.